OpenStudy (anonymous):

Paul drove from his house to work at averge speed of 40 miles per hour.The driver took him 15 minutes . If the drive home took him 20 minutes and he use the same route in rever,what is his average speed going home?

OpenStudy (dumbcow):

d = r*t d=40(1/4), 15min=1/4 hr d = 10 miles d=r*t 10=1/3 r 30 = r avg speed is 30mph

OpenStudy (anonymous):

thank